A closed tube is 10 inches long with a radius of 5 inches
find its surface area
find its volume

Answer by ankor@dixie-net.com(22740)

A closed tube is 10 inches long with a radius of 5 inches
:
find its surface area
Area of the side: circumference * length
Area of the ends: 2 * area of the circular ends
:
A = Side area + 2 ends area
A = (2*pi*5*10) + 2(pi*5^2)
A = 314.16 + 157.08
A = 471.4 cu inches
:
:
find its volume
Vol = pi*r^2*h
V = pi*5^2*10
V = 785.4 cu inches